OpenStudy (anonymous):

checking my work The length of a rectangle is 11 m and its width is 8 m. Find the area of ½ of the rectangle 22m^2

OpenStudy (anonymous):

44

Parth (parthkohli):

No Suzy, that'd make one fourth of the triangle. You have to half one of them, either length or width. Get it? :)

OpenStudy (anonymous):

half of a rectangel is right triangel by 8and 11

OpenStudy (anonymous):

yeah i got confused with the 1/2 and divided the 11 and 8 then multiplied i get thank u!

OpenStudy (lgbasallote):

\(\huge \frac{11 \times 8}{2} = \frac{88}{2} = 44\) for solution's sake haha